OpenStudy (anonymous):

Find the number of diagonals of a regular polygon whose interior angle measures 144 degrees?

ganeshie8 (ganeshie8):

what regular polygon has an interior angle measure of 144 degrees ?

ganeshie8 (ganeshie8):

exterior angle = 180-144 = 36 number of sides = 360/36 = 10 that means the polygon is a decagon, yes ?

OpenStudy (anonymous):

yep but my classmates answered 35. how they get 35 ??

ganeshie8 (ganeshie8):

use the formula http://www.mathsisfun.com/geometry/polygons-diagonals.html

OpenStudy (anonymous):

ty bro. :D

ganeshie8 (ganeshie8):

are you looking for a derivation for the n(n-3)/2 formula

ganeshie8 (ganeshie8):

?